\begin{thebibliography}{36}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Baevski et~al.(2020)Baevski, Zhou, Mohamed, and Auli]{baevski2020wav2vec}
Alexei Baevski, Yuhao Zhou, Abdelrahman Mohamed, and Michael Auli.
\newblock wav2vec 2.0: A framework for self-supervised learning of speech representations.
\newblock \emph{Advances in neural information processing systems}, 33:\penalty0 12449--12460, 2020.

\bibitem[Cabacas-Maso et~al.(2025)Cabacas-Maso, Ortega-Beltrán, Benito-Altamirano, and Ventura]{cabacasmaso2025}
Josep Cabacas-Maso, Elena Ortega-Beltrán, Ismael Benito-Altamirano, and Carles Ventura.
\newblock Enhancing facial expression recognition through dual-direction attention mixed feature networks and clip: Application to 8th abaw challenge, 2025.

\bibitem[Caron et~al.(2021)Caron, Touvron, Misra, J{\'e}gou, Mairal, Bojanowski, and Joulin]{caron2021emerging}
Mathilde Caron, Hugo Touvron, Ishan Misra, Herv{\'e} J{\'e}gou, Julien Mairal, Piotr Bojanowski, and Armand Joulin.
\newblock Emerging properties in self-supervised vision transformers.
\newblock In \emph{Proceedings of the IEEE/CVF international conference on computer vision}, pages 9650--9660, 2021.

\bibitem[Cowie et~al.(2000)Cowie, Douglas-Cowie, Savvidou, McMahon, Sawey, and Schr{\"o}der]{cowie2000feeltrace}
Roddy Cowie, Ellen Douglas-Cowie, Susie Savvidou, Edelle McMahon, Martin Sawey, and Marc Schr{\"o}der.
\newblock Feeltrace: An instrument for recording perceived emotion in real time.
\newblock In \emph{Proceedings of the ISCA Workshop on Speech and Emotion}, 2000.

\bibitem[Haque et~al.(2018)Haque, Guo, Miner, and Fei-Fei]{haque2018measuring}
Albert Haque, Michelle Guo, Adam~S Miner, and Li Fei-Fei.
\newblock Measuring depression symptom severity from spoken language and 3d facial expressions.
\newblock \emph{arXiv preprint arXiv:1811.08592}, 2018.

\bibitem[Kolhatkar et~al.(2020)Kolhatkar, Wu, Cavasso, Francis, Shukla, and Taboada]{kolhatkar2020sfu}
Varada Kolhatkar, Hanhan Wu, Luca Cavasso, Emilie Francis, Kavan Shukla, and Maite Taboada.
\newblock The sfu opinion and comments corpus: A corpus for the analysis of online news comments.
\newblock \emph{Corpus pragmatics}, 4:\penalty0 155--190, 2020.

\bibitem[Kollias(2022)]{kollias2022abaw}
Dimitrios Kollias.
\newblock Abaw: Valence-arousal estimation, expression recognition, action unit detection \& multi-task learning challenges.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 2328--2336, 2022.

\bibitem[Kollias(2023{\natexlab{a}})]{kollias2023abaw}
Dimitrios Kollias.
\newblock Abaw: learning from synthetic data \& multi-task learning challenges.
\newblock In \emph{European Conference on Computer Vision}, pages 157--172. Springer, 2023{\natexlab{a}}.

\bibitem[Kollias(2023{\natexlab{b}})]{kollias2023multi}
Dimitrios Kollias.
\newblock Multi-label compound expression recognition: C-expr database \& network.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 5589--5598, 2023{\natexlab{b}}.

\bibitem[Kollias and Zafeiriou(2019)]{kollias2019expression}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Expression, affect, action unit recognition: Aff-wild2, multi-task learning and arcface.
\newblock \emph{arXiv preprint arXiv:1910.04855}, 2019.

\bibitem[Kollias and Zafeiriou(2021{\natexlab{a}})]{kollias2021affect}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Affect analysis in-the-wild: Valence-arousal, expressions, action units and a unified framework.
\newblock \emph{arXiv preprint arXiv:2103.15792}, 2021{\natexlab{a}}.

\bibitem[Kollias and Zafeiriou(2021{\natexlab{b}})]{kollias2021analysing}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Analysing affective behavior in the second abaw2 competition.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on Computer Vision}, pages 3652--3660, 2021{\natexlab{b}}.

\bibitem[Kollias et~al.()Kollias, Schulc, Hajiyev, and Zafeiriou]{kollias2020analysing}
D Kollias, A Schulc, E Hajiyev, and S Zafeiriou.
\newblock Analysing affective behavior in the first abaw 2020 competition.
\newblock In \emph{2020 15th IEEE International Conference on Automatic Face and Gesture Recognition (FG 2020)(FG)}, pages 794--800.

\bibitem[Kollias et~al.(2019{\natexlab{a}})Kollias, Sharmanska, and Zafeiriou]{kollias2019face}
Dimitrios Kollias, Viktoriia Sharmanska, and Stefanos Zafeiriou.
\newblock Face behavior a la carte: Expressions, affect and action units in a single network.
\newblock \emph{arXiv preprint arXiv:1910.11111}, 2019{\natexlab{a}}.

\bibitem[Kollias et~al.(2019{\natexlab{b}})Kollias, Tzirakis, Nicolaou, Papaioannou, Zhao, Schuller, Kotsia, and Zafeiriou]{kollias2019deep}
Dimitrios Kollias, Panagiotis Tzirakis, Mihalis~A Nicolaou, Athanasios Papaioannou, Guoying Zhao, Bj{\"o}rn Schuller, Irene Kotsia, and Stefanos Zafeiriou.
\newblock Deep affect prediction in-the-wild: Aff-wild database and challenge, deep architectures, and beyond.
\newblock \emph{International Journal of Computer Vision}, pages 1--23, 2019{\natexlab{b}}.

\bibitem[Kollias et~al.(2021)Kollias, Sharmanska, and Zafeiriou]{kollias2021distribution}
Dimitrios Kollias, Viktoriia Sharmanska, and Stefanos Zafeiriou.
\newblock Distribution matching for heterogeneous multi-task learning: a large-scale face study.
\newblock \emph{arXiv preprint arXiv:2105.03790}, 2021.

\bibitem[Kollias et~al.(2023)Kollias, Tzirakis, Baird, Cowen, and Zafeiriou]{kollias2023abaw2}
Dimitrios Kollias, Panagiotis Tzirakis, Alice Baird, Alan Cowen, and Stefanos Zafeiriou.
\newblock Abaw: Valence-arousal estimation, expression recognition, action unit detection \& emotional reaction intensity estimation challenges.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 5888--5897, 2023.

\bibitem[Kollias et~al.(2024{\natexlab{a}})Kollias, Sharmanska, and Zafeiriou]{kollias2024distribution}
Dimitrios Kollias, Viktoriia Sharmanska, and Stefanos Zafeiriou.
\newblock Distribution matching for multi-task learning of classification tasks: a large-scale study on faces \& beyond.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pages 2813--2821, 2024{\natexlab{a}}.

\bibitem[Kollias et~al.(2024{\natexlab{b}})Kollias, Tzirakis, Cowen, Zafeiriou, Kotsia, Baird, Gagne, Shao, and Hu]{kollias20246th}
Dimitrios Kollias, Panagiotis Tzirakis, Alan Cowen, Stefanos Zafeiriou, Irene Kotsia, Alice Baird, Chris Gagne, Chunchang Shao, and Guanyu Hu.
\newblock The 6th affective behavior analysis in-the-wild (abaw) competition.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 4587--4598, 2024{\natexlab{b}}.

\bibitem[Kollias et~al.(2024{\natexlab{c}})Kollias, Zafeiriou, Kotsia, Dhall, Ghosh, Shao, and Hu]{kollias20247th}
Dimitrios Kollias, Stefanos Zafeiriou, Irene Kotsia, Abhinav Dhall, Shreya Ghosh, Chunchang Shao, and Guanyu Hu.
\newblock 7th abaw competition: Multi-task learning and compound expression recognition.
\newblock \emph{arXiv preprint arXiv:2407.03835}, 2024{\natexlab{c}}.

\bibitem[Kollias et~al.(2025)Kollias, Tzirakis, Cowen, Zafeiriou, Kotsia, Granger, Pedersoli, Bacon, Baird, Gagne, et~al.]{kollias2025advancements}
Dimitrios Kollias, Panagiotis Tzirakis, Alan Cowen, S Zafeiriou, I Kotsia, Eric Granger, Marco Pedersoli, Simon Bacon, Alice Baird, C Gagne, et~al.
\newblock Advancements in affective and behavior analysis: The 8th abaw workshop and competition.
\newblock 2025.

\bibitem[Lawrence and Lin(1989)]{lawrence1989concordance}
I Lawrence and Kuei Lin.
\newblock A concordance correlation coefficient to evaluate reproducibility.
\newblock \emph{Biometrics}, pages 255--268, 1989.

\bibitem[Mollahosseini et~al.(2017)Mollahosseini, Hasani, and Mahoor]{mollahosseini2017affectnet}
Ali Mollahosseini, Behzad Hasani, and Mohammad~H Mahoor.
\newblock Affectnet: A database for facial expression, valence, and arousal computing in the wild.
\newblock \emph{IEEE Transactions on Affective Computing}, 10\penalty0 (1):\penalty0 18--31, 2017.

\bibitem[Qin et~al.(2024)Qin, Leichner, Delakis, Fornoni, Luo, Yang, Wang, Banbury, Ye, Akin, et~al.]{qin2024mobilenetv4}
Danfeng Qin, Chas Leichner, Manolis Delakis, Marco Fornoni, Shixin Luo, Fan Yang, Weijun Wang, Colby Banbury, Chengxi Ye, Berkin Akin, et~al.
\newblock Mobilenetv4: universal models for the mobile ecosystem.
\newblock In \emph{European Conference on Computer Vision}, pages 78--96. Springer, 2024.

\bibitem[Savchenko(2024)]{savchenko2024hsemotion}
Andrey~V Savchenko.
\newblock Hsemotion team at the 6th abaw competition: Facial expressions, valence-arousal and emotion intensity prediction.
\newblock \emph{arXiv preprint arXiv:2403.11590}, 2024.

\bibitem[Savchenko(2025)]{savchenko2025hsemotion}
Andrey~V Savchenko.
\newblock Hsemotion team at abaw-8 competition: Audiovisual ambivalence/hesitancy, emotional mimicry intensity and facial expression recognition.
\newblock \emph{arXiv preprint arXiv:2503.10399}, 2025.

\bibitem[Sikander and Anwar(2018)]{sikander2018driver}
Gulbadan Sikander and Shahzad Anwar.
\newblock Driver fatigue detection systems: A review.
\newblock \emph{IEEE Transactions on Intelligent Transportation Systems}, 20\penalty0 (6):\penalty0 2339--2352, 2018.

\bibitem[Tolstikhin et~al.(2021)Tolstikhin, Houlsby, Kolesnikov, Beyer, Zhai, Unterthiner, Yung, Steiner, Keysers, Uszkoreit, et~al.]{tolstikhin2021mlp}
Ilya~O Tolstikhin, Neil Houlsby, Alexander Kolesnikov, Lucas Beyer, Xiaohua Zhai, Thomas Unterthiner, Jessica Yung, Andreas Steiner, Daniel Keysers, Jakob Uszkoreit, et~al.
\newblock Mlp-mixer: An all-mlp architecture for vision.
\newblock \emph{Advances in neural information processing systems}, 34:\penalty0 24261--24272, 2021.

\bibitem[Tzirakis et~al.(2017)Tzirakis, Trigeorgis, Nicolaou, Schuller, and Zafeiriou]{tzirakis2017end}
Panagiotis Tzirakis, George Trigeorgis, Mihalis~A Nicolaou, Bj{\"o}rn~W Schuller, and Stefanos Zafeiriou.
\newblock End-to-end multimodal emotion recognition using deep neural networks.
\newblock \emph{IEEE Journal of selected topics in signal processing}, 11\penalty0 (8):\penalty0 1301--1309, 2017.

\bibitem[Tzirakis et~al.(2018{\natexlab{a}})Tzirakis, Zafeiriou, and Schuller]{tzirakis2018end2you}
Panagiotis Tzirakis, Stefanos Zafeiriou, and Bjorn~W Schuller.
\newblock End2you--the imperial toolkit for multimodal profiling by end-to-end learning.
\newblock \emph{arXiv preprint arXiv:1802.01115}, 2018{\natexlab{a}}.

\bibitem[Tzirakis et~al.(2018{\natexlab{b}})Tzirakis, Zhang, and Schuller]{tzirakis2018end}
Panagiotis Tzirakis, Jiehao Zhang, and Bjorn~W Schuller.
\newblock End-to-end speech emotion recognition using deep neural networks.
\newblock In \emph{2018 IEEE international conference on acoustics, speech and signal processing (ICASSP)}, pages 5089--5093. IEEE, 2018{\natexlab{b}}.

\bibitem[Tzirakis et~al.(2021{\natexlab{a}})Tzirakis, Chen, Zafeiriou, and Schuller]{tzirakis2021end}
Panagiotis Tzirakis, Jiaxin Chen, Stefanos Zafeiriou, and Bj{\"o}rn Schuller.
\newblock End-to-end multimodal affect recognition in real-world environments.
\newblock \emph{Information Fusion}, 68:\penalty0 46--53, 2021{\natexlab{a}}.

\bibitem[Tzirakis et~al.(2021{\natexlab{b}})Tzirakis, Nguyen, Zafeiriou, and Schuller]{tzirakis2021speech}
Panagiotis Tzirakis, Anh Nguyen, Stefanos Zafeiriou, and Bj{\"o}rn~W Schuller.
\newblock Speech emotion recognition using semantic information.
\newblock In \emph{ICASSP 2021-2021 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP)}, pages 6279--6283. IEEE, 2021{\natexlab{b}}.

\bibitem[Vu et~al.(2023)Vu, Huynh, Nguyen, and Kim]{vu2023ensemble}
Ngoc~Tu Vu, Van~Thong Huynh, Trong~Nghia Nguyen, and Soo-Hyung Kim.
\newblock Ensemble spatial and temporal vision transformer for action units detection.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 5770--5776, 2023.

\bibitem[Zafeiriou et~al.(2017)Zafeiriou, Kollias, Nicolaou, Papaioannou, Zhao, and Kotsia]{zafeiriou2017aff}
Stefanos Zafeiriou, Dimitrios Kollias, Mihalis~A Nicolaou, Athanasios Papaioannou, Guoying Zhao, and Irene Kotsia.
\newblock Aff-wild: Valence and arousal ‘in-the-wild’challenge.
\newblock In \emph{Computer Vision and Pattern Recognition Workshops (CVPRW), 2017 IEEE Conference on}, pages 1980--1987. IEEE, 2017.

\bibitem[Zhang et~al.(2016)Zhang, Zhang, Li, and Qiao]{zhang2016joint}
Kaipeng Zhang, Zhanpeng Zhang, Zhifeng Li, and Yu Qiao.
\newblock Joint face detection and alignment using multitask cascaded convolutional networks.
\newblock \emph{IEEE signal processing letters}, 23\penalty0 (10):\penalty0 1499--1503, 2016.

\end{thebibliography}

\documentclass[10pt,twocolumn,letterpaper]{article}

\usepackage{cvpr}              

\definecolor{cvprblue}{rgb}{0.21,0.49,0.74}
\usepackage[pagebackref,breaklinks,colorlinks,allcolors=cvprblue]{hyperref}

\def\paperID{*****} 
\def\confName{CVPR}
\def\confYear{2025}

\title{Lightweight Models for Emotional Analysis in Video}

\author{Quoc-Tien Nguyen, Hong-Hai Nguyen, Van-Thong Huynh\thanks{Corresponding author}\\
Dept. of ITS, FPT University\\
Ho Chi Minh City, 71216, Vietnam\\
{\tt\small \{tiennq27,hainh51,thonghv4\}@fe.edu.vn}
}

\begin{document}
\maketitle
\begin{abstract}


In this study, we present an approach for efficient spatiotemporal feature extraction using MobileNetV4 and a multi-scale 3D MLP-Mixer-based temporal aggregation module. MobileNetV4, with its Universal Inverted Bottleneck (UIB) blocks, serves as the backbone for extracting hierarchical feature representations from input image sequences, ensuring both computational efficiency and rich semantic encoding. To capture temporal dependencies, we introduce a three-level MLP-Mixer module, which processes spatial features at multiple resolutions while maintaining structural integrity. Experimental results on the ABAW 8th competition demonstrate the effectiveness of our approach, showing promising performance in affective behavior analysis. By integrating an efficient vision backbone with a structured temporal modeling mechanism, the proposed framework achieves a balance between computational efficiency and predictive accuracy, making it well-suited for real-time applications in mobile and embedded computing environments.
\end{abstract}    
\section{Introduction}
\label{sec:intro}

Human emotion recognition is a subfield of human behavior analysis that has got significant interest from researchers in artificial intelligence, psychology, and human-computer interaction. By utilizing various types of data, such as audio, images, and text, researchers can analyze and predict human emotions as well as continuous actions. Advancements in deep learning has significantly improved the accuracy and efficiency of emotion recognition systems. These technologies can detect emotions in real time, allowing applications across a wide range of fields. Understanding human emotions and behavior can create various applications in areas such as healthcare, autonomous vehicles, robotics, and more \cite{kollias2022abaw, haque2018measuring, sikander2018driver, kolhatkar2020sfu}. For example, emotion recognition is increasingly being used in mental health monitoring, where it helps identify signs of depression or stress based on behavioral patterns. In marketing, consumer emotion recognition enables brands to tailor their products and advertising strategies to evoke desired emotional responses. Moreover, in human-computer interaction, emotion-aware systems are enhancing user experiences by adapting to users’ emotional state.

The Emotion recognition has many benefits in life, but emotion recognition still faces various challenges in real-world applications. Human emotions are complex and influenced by various factors such as gender, age, and context, etc. Therefore, the 8th Affective \& Behavior Analysis in-the-Wild (ABAW8) workshop \cite{kollias2025advancements} specifically addresses the complex challenges inherent in analyzing human affective states and behavioral patterns in unconstrained, real-world scenarios. Unlike controlled laboratory environments, in-the-wild settings present numerous variables including diverse lighting conditions, varying head poses, occlusions, and spontaneous expressions that significantly complicate accurate analysis. 

The challenge includes six tasks: Valence-Arousal (VA) Estimation, Expression (EXPR) Classification, Action Unit(AU) Detection, Compound Expression (CE) Recognition, Emotional Mimicry Intensity (EMI) Estimation, Ambivalence/Hesitancy (AH) Recognition. The challenges utilize the Aff-Wild2 \cite{zafeiriou2017aff,kollias20247th,kollias20246th,kollias2024distribution,kollias2023abaw2,kollias2023multi,kollias2023abaw,kollias2022abaw,kollias2021analysing,kollias2021affect,kollias2021distribution,kollias2020analysing,kollias2019expression,kollias2019deep,kollias2019face}, C-EXPR-DB \cite{kollias2023multi}, HUME-Vidmimic2, and BAH datasets, providing a comprehensive evaluation framework for affective behavior analysis methodologies. In this paper, we solve the tasks such as Valence-Arousal (VA) Estimation, Action Unit(AU) Detection, Emotional Mimicry Intensity (EMI) Estimation, Ambivalence/Hesitancy (AH) Recognition. The VA task predicts valence and arousal in video sequences, while the AU task detects the presence of 12 action units (AUs) in each video frame. The EMI task estimates the intensity of six emotional dimensions (Admiration, Amusement, Determination, Empathic Pain, Excitement, and Joy). Lastly, the AH task identifies the presence or absence of ambivalence or hesitancy in each frame.




\section{Method}\label{sec:method}

\subsection{Visual feature extraction}

The visual feature extraction process is facilitated by a small variant of MobileNetV4~\cite{qin2024mobilenetv4}, pretrained on AffectNet dataset~\cite{mollahosseini2017affectnet}. MobileNetV4~\cite{qin2024mobilenetv4} introduces a universally efficient architecture optimized for mobile devices, integrating the Universal Inverted Bottleneck (UIB) and Mobile Multi-Query Attention (Mobile MQA) blocks to enhance feature extraction efficiency. The UIB block unifies key architectural components, including Inverted Bottleneck (IB), ConvNext, Feed Forward Network (FFN), and an Extra Depthwise (ExtraDW) variant, providing flexibility in spatial and channel mixing while improving computational efficiency. Mobile MQA accelerates attention mechanisms by over 39\% on mobile accelerators, further optimizing feature representation. Additionally, an advanced neural architecture search (NAS) strategy refines MobileNetV4’s design, ensuring mostly Pareto-optimal performance across CPUs, DSPs, GPUs, and dedicated accelerators like Google's EdgeTPU. 
In this study, we fed a sequence of $224\times224\times3$ images into the MobileNetV4 to extract multi-scale feature maps from input frames. The backbone is configured to output hierarchical feature representations at different spatial resolutions, allowing for rich semantic extraction. To maintain efficiency while preserving important spatial details, all but the final layers of the backbone are frozen during training. The extracted feature maps serve as input to the subsequent temporal modeling module.

\subsection{Temporal aggregation module}
To model temporal dependencies in the extracted features, we incorporate a multiscale 3D MLP-Mixer-based~\cite{tolstikhin2021mlp} to build temporal aggregation module (TAM). This module processes the sequential feature maps using multiple levels of spatial granularity. TAM consist of three mixer layers operate on different feature resolutions: (1) a high-resolution mixer for $28\times28$ feature maps, (2) a mid-resolution mixer with input size of $14\times14$, and (3) a low-resolution mixer for highly detail feature maps of size $7\times7$. Each mixer employs 3D MLP-based transformations to capture temporal relationships while preserving spatial structure. Finally, a fully connected layer maps the aggregated feature representations to the target output space, ensuring effective sequence-level prediction.

\section{Experiments and Results}\label{sec:exp_result}
\subsection{Dataset}
This subsection provides a brief summary of the datasets used in each challenge, such as Action Unit (AU) Detection, Valence-Arousal (VA) Estimation, Emotional Mimicry Intensity (EMI) Estimation.
\paragraph{Action Unit (AU) Detection}
This challenge uses a data set that includes 542 videos with annotations for 12 Action Units (AU), which represent facial muscle movements, including brow raisers, cheek raisers, lip tighteners, and jaw drops. The data set consists of 2,627,632 frames captured from 438 unique subjects. Annotations were developed through a semiautomatic methodology that integrates both manual and computational techniques. The data set has been divided into three subsets: a training set (295 videos), a validation set (105 videos), and a testing set (142 videos). \autoref{tab:au_dis} provides the distribution of the AU annotations of the dataset.


\paragraph{Valence-Arousal (VA) Estimation}
This challenge utilizes an expanded version of the Aff-Wild2 database, including 594 videos annotated for valence and arousal. The data set consists of 2,993,081 frames captured from 584 subjects. Significantly, sixteen videos contain dual subjects, both of whom received independent annotations. Four expert annotators evaluated the data set following the methodology detailed in \cite{cowie2000feeltrace}, continuous valence, and arousal values within the range of [-1, 1].

To maintain subject independence across experimental protocols, the data set has been partitioned into three discrete subsets: a training set (356 videos), a validation set (76 videos), and a testing set (162 videos). This division ensures that individual subjects appear exclusively in one subset.
\paragraph{Emotional Mimicry Intensity (EMI) Estimation.}

 The EMI Challenge considers emotional mimicry through the HUME-Vidmimic2 dataset, which includes more than 30 hours of audiovisual recordings from 557 participants. This data set was collected in naturalistic environments where participants used their webcams to mimic facial and vocal expressions presented in seed videos, subsequently self-evaluating their mimicry performance on a scale of 0-100.
The data set has been systematically partitioned according to the distribution detailed in \autoref{tab:HUME_stat}, which presents comprehensive statistics for each subset. To facilitate analysis, participants are provided with facial detections extracted from videos using MTCNN \cite{zhang2016joint}, processed at a sampling rate of 6 frames per second.

\begin{table}
    \centering
    \caption{HUME-Vidmimic2 partition statistics.
    \label{tab:HUME_stat}}
    \begin{tabular}{c c c} \toprule
        Partition &  Duration & \#Samples \\ 
        & (HH:MM:SS) & \\ \midrule
        Train & 15:07:03  & 8072 \\
        Validation & 9:12:02 & 4588  \\
        Test & 9:04:05  & 4586 \\
        \bottomrule
    \end{tabular}
\end{table}

Furthermore, to enable the development of end-to-end methodological approaches \cite{tzirakis2017end,tzirakis2018end2you,tzirakis2018end,tzirakis2021end,tzirakis2021speech}, participants receive pre-extracted features derived from the raw audiovisual data, specifically: facial features processed using Vision Transformer (ViT) \cite{caron2021emerging}, audio signals processed using Wav2Vec 2.0 \cite{baevski2020wav2vec}.

\paragraph{ Ambivalence/Hesitancy Recognition}


\paragraph{Action Unit (AU) Detection}
The $F1$ score uses precision and recall to calculate which ensure a robust assessment of classification performance. The equation of the $F_1$ score is as follows:

 \begin{equation}
    \begin{split}
        \label{eqn:F1_score}
        {F_1} &= \frac{2*precision*recall}{precision+recall}
    \end{split}
\end{equation}

The average F1 score is used to evaluate 12 AUs. The F1 score ranges from 0 to 1, where 1 represents perfect and 0 represents the worst performance. The formula is expressed as follows:
 \begin{equation}
    \begin{split}
        F_{\mathrm{AU}} &= \frac{\sum_{au}F_{1}^{au}}{12}
    \end{split}
\end{equation}

\paragraph{Valence-Arousal (VA) Estimation} In this task, the Concordance Correlation Coefficient (CCC) \cite{lawrence1989concordance}, $\mathcal{P}$ is used to evaluate the performance of the model. The CCC is a measure that is used to evaluate the agreement between two continuous variables. The CCC range is from [-1,1] where -1 is a negative correlation, 0 is no correlation, and 1 is a high correlation. The formula is expressed as follows:

 \begin{equation}
    \begin{split}
        \label{eqn:va_combined}
        \mathcal{P}_{\mathrm{VA}} &= \frac{\mathcal{P}_\mathrm{V} + \mathcal{P}_{\mathrm{A}}}{2}
    \end{split}
\end{equation}
where $\mathcal{P}_\mathrm{V}$ and $\mathcal{P}_\mathrm{A}$ are the CCC of valence and arousal, respectively, which is defined as
\begin{equation}
    \begin{split}
        \label{eqn:ccc}
        \mathcal{P} &= \frac{2 \rho \sigma_{\hat Y} \sigma_{Y} }{\sigma_{\hat Y}^2 + \sigma_{Y}^2 + (\mu_{\hat Y} - \mu_{Y})^2}
    \end{split}
\end{equation}

where $\mu_{Y}$ was the mean of the label $Y$, $\mu_{\hat Y}$ was the mean of prediction $\hat Y$, $\sigma_{\hat Y}$ and $\sigma_{Y}$ were the corresponding standard deviations,  $\rho$ was the Pearson correlation coefficient between $\hat Y$and $Y$. 

\paragraph{Emotional Mimicry Intensity (EMI) Estimation.}
The average Pearson’s Correlation Coefficient ($\rho$)is used to measure the six emotion dimensions:
 \begin{equation}
    \begin{split}
        \mathcal{P}_{\mathrm{EMI}} &= \frac{\sum_{i=1}^{6}{\rho}^{i}}{6}
    \end{split}
\end{equation}



\autoref{tab:au_task} shown results of our approach compared to previous studies on Affwild2 dataset~\cite{kollias2021affect}.
\begin{table}[htbp]
    \centering
    \caption{Evaluation on Action Unit detection task with F1 score.}
    \label{tab:au_task}
    \begin{tabular}{l c c}\toprule
       Methods  & Validation & Test \\ \midrule
       Regnet-ViT~\cite{vu2023ensemble} & 0.5280 & - \\ 
       EmotiEffNet~\cite{savchenko2024hsemotion,savchenko2025hsemotion} & 0.537 & 0.488 \\ 
       AIWELL-UOC~\cite{cabacasmaso2025} & 0.451 & 0.417 \\ \midrule
         Ours with 3 level mixer & 0.5369 & - \\
         Ours with 2 level mixer & 0.5338 & - \\
         Ours - single mixer & 0.5441 & 0.484\\
         \bottomrule
    \end{tabular}
\end{table}

{
    \small
    \bibliographystyle{ieeenat_fullname}
    \bibliography{main}
}


\end{document}